Circular No. 7963

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002eu D. Weisz and W. D. Li, University of California at Berkeley, report

the LOTOSS (cf. IAUC 7906) discovery of an apparent supernova on an unfiltered KAIT CCD image taken on Aug. 30.4 UT (mag ∼ 17.4). The new object was confirmed on an earlier image taken on Aug. 25.4 (mag ∼ 18.8). SN 2002eu is located at α = 1^{h}49^{m}43^{s}.60, δ = +32^{o}37′43′′.0 (J2000.0), which is 10′′.5 east and 11′′.5 north of the nucleus of an anonymous galaxy. A KAIT image taken on Aug. 20.4 showed nothing at this position (limiting mag ∼ 19.0).

SUPERNOVA 2002ej IN NGC 6324 S. Desidera, E. Giro, A. Della Valle, S. Benetti, G. Altavilla, A. Pas-

torello, and M. Turatto, Osservatorio Astronomico di Padova; and E. Cappellaro, Osservatorio Astronomico di Capodimonte, write: “A noisy spectrum (range 400–780 nm, resolution 2.5 nm) of SN 2002ej (cf. IAUC 7951), obtained on Aug. 30.0 UT with the Asiago 1.82-m telescope (+ AFOSC), shows that it is a type-II supernova, ∼ two weeks after maximum brightness. The spectrum is dominated by a broad Hα line with a P-Cyg profile, from which an expansion velocity of ∼ 8100 km/s is deduced (assuming the NED recession velocity of 4866 km/s for the parent galaxy).”

COMETS C/2002 Q6, C/2002 Q7, C/2002 Q8 (SOHO) Further to IAUC 7956, D. Hammer reports measurements for three

additional comets found by X. Leprette (C/2002 Q6), J. Sachs (C/2002 Q7), and X.-M. Zhou (C/2002 Q8) on SOHO website C2 images; C/2002 Q7 was also seen in C3 coronagraph images. C/2002 Q7 is a Kreutz sungrazer, and C/2002 Q8 appears to be a member of the Kracht group (cf. IAUC 7952). The reduced astrometry and orbital elements by B. G. Marsden appear on MPEC 2002-Q46.

Comet  2002 UT  α_{2000}  δ_{2000}
C/2002 Q6 Aug. 23.451  10^{h}00.^{m}1  +11^{o}17′
C/2002 Q7  23.696  9 41.1  + 9 49
C/2002 Q8  26.018  10 11.6  +11 44
2002 August 31  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green